Solve the problem and express the solution in a + bi form
(3 - sqrt(-5))*(1 + sqrt(-1))

Answer by dabanfield(803)

(3 - sqrt(-5))*(1 + sqrt(-1)) =
3*1 + 3*sqrt(-1) - 1*sqrt(-5) - sqrt(-5*-1)
3 + 3*sqrt(-1) - sqrt(-1*5) - sqrt(5) =
3 + 3*i - i*sqrt(5) - sqrt(5) =
[3-sqrt(5)] + [3-sqrt(5)]*i